# Prototype-Based Disentanglement for Controllable Dysarthric Speech Synthesis

### v1

@@ Abstract
Dysarthric speech exhibits high variability and limited labeled data, posing major challenges for both automatic speech recognition (ASR) and assistive speech technologies. Existing approaches rely on synthetic data augmentation or speech reconstruction, yet often entangle speaker identity with pathological articulation, limiting controllability and robustness.
In this paper, we propose ProtoDisent-TTS, a prototype-based disentanglement TTS framework built on a pre-trained text-to-speech backbone that factorizes speaker timbre and dysarthric articulation within a unified latent space. A pathology prototype codebook provides interpretable and controllable representations of healthy and dysarthric speech patterns, while a dual-classifier objective with a gradient reversal layer enforces invariance of speaker embeddings to pathological attributes. Experiments on the TORGO dataset demonstrate that this design enables bidirectional transformation between healthy and dysarthric speech, leading to consistent ASR performance gains and robust, speaker-aware speech reconstruction. [^1] Code and audio samples: https://mors20.github.io/ProtoDisent-TTS/
@@ 1 Introduction
Dysarthria is a motor speech disorder caused by impaired control or weakness of the muscles involved in speech production [1]. It often results in slurred, slow, or imprecise articulation that hinders communication and poses significant challenges for automatic speech recognition (ASR) and speech-based assistive technologies, due to the high variability of pathological speech and the limited availability of labeled data [2, 3]. Prior work primarily follows two main research directions.
@@ 2.1 Problem formulation and framework overview
$\mathcal{P}=\{\mathbf{p}_{0},\mathbf{p}_{1},\dots,\mathbf{p}_{n}\},\quad\mathbf{p}_{i}\in\mathbb{R}^{D}$ (1)
, according to the selected prototype index $k$.

### v2

@@ Abstract
Dysarthric speech exhibits high variability and limited labeled data (accurately transcribed recordings), posing major challenges for both automatic speech recognition (ASR) and assistive speech technologies. Current methods depend on synthetic data or reconstructed speech, yet they often fail to separate speaker identity from pathological speech patterns, thereby reducing controllability and robustness.
In this paper, we propose ProtoDisent-TTS, a prototype-based disentanglement TTS framework built on a pre-trained text-to-speech backbone that factorizes speaker timbre and dysarthric articulation within a unified latent space. A pathology prototype codebook (a defined set of speech patterns) provides interpretable and controllable representations of healthy and dysarthric speech, while a dual-classifier objective with a gradient reversal layer ensures speaker identity remains separate from these pathological attributes. Experiments on the TORGO dataset demonstrate that our design can bidirectionally convert healthy and dysarthric speech, resulting in consistent ASR performance gains and robust, speaker-aware speech reconstruction. [^1] Code and audio samples: https://mors20.github.io/ProtoDisent-TTS. The paper has been submitted to Interspeech 2026.
@@ 1 Introduction
Dysarthria is a neurological motor speech disorder primarily caused by damage to the central or peripheral nervous system that impairs control over the muscles used for speech production [1]. It often results in slurred, slow, or imprecise articulation that hinders communication and poses significant challenges for automatic speech recognition (ASR) and speech-based assistive technologies, due to the high variability of pathological speech and the limited availability of labeled data [2, 3]. Prior work primarily follows two main research directions.
@@ 2.1 Problem formulation and framework overview
$\mathcal{P}=\{\mathbf{p}_{0},\mathbf{p}_{1},\dots,\mathbf{p}_{n}\},\quad\mathbf{p}_{i}\in\mathbb{R}^{D},$ (1)
according to the selected prototype index $k$.